# Neurodynamics-Driven Coupled Neural P Systems for Multi-Focus Image Fusion

### v2

@@ Abstract
Multi-focus image fusion (MFIF) is a crucial technique in image processing, with a key challenge being the generation of decision maps with precise boundaries. However, traditional methods based on heuristic rules and deep learning methods with black-box mechanisms are difficult to generate high-quality decision maps. To overcome this challenge, we introduce neurodynamics-driven coupled neural P (CNP) systems, which are third-generation neural computation models inspired by spiking mechanisms, to enhance the accuracy of decision maps. Specifically, we first conduct an in-depth analysis of the model’s neurodynamics to identify the constraints between the network parameters and the input signals. This solid analysis avoids abnormal continuous firing of neurons and ensures the model accurately distinguishes between focused and unfocused regions, generating high-quality decision maps for MFIF. Based on this analysis, we propose a Neurodynamics-Driven CNP Fusion model (ND-CNPFuse) tailored for the challenging MFIF task. Unlike current ideas of decision map generation, ND-CNPFuse distinguishes between focused and unfocused regions by mapping the source image into interpretable spike matrices. By comparing the number of spikes, an accurate decision map can be generated directly without any post-processing. Extensive experimental results show that ND-CNPFuse achieves new state-of-the-art performance on four classical MFIF datasets, including Lytro, MFFW, MFI-WHU, and Real-MFF. The code is available at https://github.com/MorvanLi/ND-CNPFuse.
@@ 1 Introduction
Multi-focus image fusion (MFIF) [40] is an important task in computer vision that aims to generate an all-in-focus image by fusing multiple images of the same scene captured at varying focal depths. The fused image improves both perceptual quality and information richness, thereby providing a reliable input for a range of downstream vision applications such as action detection [28] and object detection [13]. By integrating the focused regions from source images with different focus settings, MFIF effectively alleviates the blurring artifacts introduced by the limited depth of field inherent in imaging systems.
@@
Figure 1: Workflow comparison between the existing DL-based decision method and the proposed ND-CNPFuse. ND-CNPFuse encodes source images into spike matrices and generates the decision map by comparing spike counts. Focused regions generate more spikes than unfocused ones, consistent with human visual perception [9].
@@
Before the advent of deep learning (DL), MFIF methods were performed in the transform [14] and spatial domains [41]. The effectiveness of these methods often relies heavily on manually designed fusion rules and strong prior assumptions, which may not always be applicable or sufficiently accurate.
@@
In recent years, DL-based approaches have become the dominant paradigm in MFIF, achieving remarkable progress [16]. Some studies have applied DL techniques to end-to-end image fusion frameworks [35, 33, 26]. Different end-to-end methods may use various network structures, but they generally consist of an encoder network for extracting image features, a fusion network for merging features, and a decoder network for reconstructing the fused features. Another line of studies draws inspiration from spatial domain methods, suggesting the use of DL techniques to generate decision maps [19, 29, 24, 43, 7]. This process involves training a network to map the source image to a focus map, ultimately producing decision maps. Compared to end-to-end methods, decision-based methods directly retain information from the source image at the pixel level, avoiding the spatial consistency issues of end-to-end methods. Therefore, more research has been done on DL-based decision map methods. However, DL-based decision map methods still have evident drawbacks. The most crucial one is that the internal working mechanism is difficult to explain, hindering a clear understanding of the mapping between source images and the decision map [30]. This will result in nuisance false edges and burrs in the generated decision map, as can be intuitively observed in Fig. 1(a).
@@
Coupled neural P (CNP) systems [22] are a class of neural-like computing models inspired by the mechanism of spiking neurons, and are classified as third-generation neural networks. A CNP system generates synchronous spikes through the mutual coupling of neurons, where synchronous pulses of neurons are treated as 0 or 1. After multiple iterations, the system produces a sequence of binary images that record the firing states of each neuron at every time step. By comparing the spike counts of neurons, the degree of focus corresponding to each pixel can be effectively estimated, enabling the differentiation between focused and defocused regions. This is particularly well-suited for decision map generation in the MFIF task, which serves as the core motivation for introducing CNP systems into MFIF. However, when CNP systems are directly applied to MFIF, neurons may suffer abnormal continuous firing. In this case, the spike counts cannot accurately reflect the focus differences of the source images, which affects the quality of the decision map. This issue stems from an inappropriate constraint between network parameters and input signals.
@@
To address this problem, we first conduct a detailed study on the neurodynamics of CNP neurons. This analysis reveals the underlying constraints between network parameters and input signals, thereby preventing abnormal continuous firing in neurons. A solid understanding of the neural firing mechanism is helpful for setting appropriate parameters and provides a theoretical basis for effectively applying CNP systems in MFIF. Building on the neurodynamics-driven CNP system, we then propose a new fusion method called ND-CNPFuse. As illustrated in Fig. 1(b), ND-CNPFuse provides an interpretable and accurate decision map generation workflow, in contrast to existing DL-based MFIF methods that rely on black-box mechanisms (Fig. 1(a)).
@@ 2.1 DL-based MFIF methods
End-to-end methods. These methods utilize the powerful fitting capabilities of deep neural networks to directly generate fused images by learning features from the source images. Representative approaches include those based on convolutional neural networks [42, 32, 36, 3], generative adversarial networks [37, 6], and Transformer architectures [20, 44, 27]. It is challenging for these methods to maintain the spatial consistency of the source images.
Decision-map methods. These methods formulate MFIF as a pixel-level classification task by training a deep neural network to predict decision maps for the focused and defocused regions. Liu et al. [15] pioneered the application of this technique in the MFIF task, but the accuracy of the generated decision maps is still limited. To improve the accuracy of decision maps, subsequent studies have introduced ensemble learning [1], attention mechanisms [31], regression pair learning [12], multi-scale architectures [17], edge preservation techniques [30], and explicit defocus blur modelling [25]. Due to the unclear working mechanism of the network, there are interfering pseudo-edges and burrs in the decision maps. How to generate high-quality decision maps for the MFIF task remains to be explored.
@@ 3.1 Mathematical model of CNP systems
$T_{ij}(t+1)=\left\{\begin{array}[]{ll}T_{ij}(t)-\tau+\lambda P_{ij}(t),&\text{if}~\sigma_{ij}~\text{fires}\\ T_{ij}(t),&\text{otherwise}\end{array}\right$. (5)
where subscripts $\left(i,j\right)$ denote the coordinates of the neuron, $\left(k,l\right)$ denote the coordinates of its neighboring neurons, and $t$ represents the current iteration count. $u,v,\tau$ indicate the spikes consumed by three units. $I_{ij}(t)$ is an external input. $W_{kl}$ characterizes synaptic weight. $P_{kl}$ symbolizes spikes received from the surrounding neurons $\sigma_{kl}$, where neighborhood radius is $r$ ($\sigma_{kl}\in\delta_{r}$). $\lambda$ is a threshold weight.
@@ 3.2 Why neurodynamics analysis is necessary?
In certain cases, CNP neurons may enter a state of continuous firing, generating spikes at each time step. As illustrated in Fig. 2(c) and Fig. 2(d). This uncontrolled spiking behavior can lead to model failure, as the output no longer reflects the features of the input signal. For the MFIF task, such outputs do not effectively distinguish between focused and defocused regions, thereby affecting the generation of accurate decision maps. In the following, we analyze this invalid state and identify parameter configurations that can effectively prevent it.
Analysis of continuous firing condition. We analyze the continuous firing condition of CNP neurons to derive constraints between parameters and the input signal. Suppose $t$ tends to infinity, all neurons fire. In this case, the state update equation becomes:
$U(t+1)$ $=U(t)-u+I(t)+\sum\limits_{\sigma_{kl}\in\delta_{r}}W_{kl}P_{kl}(t),$ (6)
$V(t+1)$ $=V(t)-v+\sum\limits_{\sigma_{kl}\in\delta_{r}}W_{kl}P_{kl}(t),$ (7)
$T(t+1)$ $=T(t)-\tau+\lambda P(t).$ (8)
According to the theory in [22], the spike consumption of the three units in CNP neurons exhibits a linear relationship. Thus, the consumption rules can be described as:
@@
$U(t)$ $=I\frac{1-\alpha^{t}}{1-\alpha}+\sum_{i=0}^{t-1}K(i)\alpha^{t-i-1}.$ (10)
@@
Let us clarify the initial conditions: $U(0)$ and $K(0)$ are both zero. Define $K(i)=\sum_{\sigma_{kl}\in\delta_{r}}W_{kl}P_{kl}(t-1)$. Next, we describe $U(1),U(2),U(3),\ldots,U(t)$ step by step:
@@
$U(t)$ $=I\frac{1-\alpha^{t}}{1-\alpha}+\sum_{i=0}^{t-1}K(i)\alpha^{t-i-1}.$
@@
Figure 3: Overview of ND-CNPFuse for MFIF. Fused image F quality primarily relies on the decision map generated by the ND-CNP system. A and B are the multi-focus source images. $\otimes$ and $\oplus$ denote element-wise multiplication and addition.
@@
$\begin{split}&\biggl(I\frac{1-\alpha^{t}}{1-\alpha}+\sum_{i=0}^{t-1}K(i)\alpha^{t-i-1}\biggl)\biggl(1+\sum_{i=0}^{t-1}K(i)\beta^{t-i-1}\biggl)\\ &\quad\quad\quad\quad\quad\quad\quad\quad>\lambda\frac{1-\gamma^{t-1}}{1-\gamma}.\end{split}$ (17)
Given continuous firing, $K(i)$ is only related to the weights between neurons. At this point, $K(i)$ can be written as $\text{sum}(W)$ (constant synaptic input).
Eq. (17) can be expressed as:
@@
When $t$ goes to infinity, we obtain Eq. (19)
@@
According to Theorem 4, the external input to a CNP neuron must remain below its internal state to ensure proper functionality, as demonstrated below
$I\leq\frac{\lambda(1-\alpha)(1-\beta)}{(1-\gamma)(1-\beta+\operatorname{sum}(W)))}-\operatorname{sum}(W).$ (21)
Corollary 1 reveals the necessary constraints between parameters and input signals for the proper functioning of CNP neurons. Detailed CNP neurons parameter configurations (e.g., $u$, $v$, $\tau$, $\lambda$, and W) derived from Corollary 1 are provided in Appendix A.2. Noting that the internal parameters of neurons are automatically configured based on input MFIF images, eliminating the manual tuning, and demonstrating transferability and generalization across different datasets. This helps to improve the ability of CNP systems to distinguish between focused and defocused regions in MFIF tasks, thereby improving the quality of the generated decision map.
@@ 4.2 Overview of the architecture
MGDN 0.7393 0.4558 0.8505 26.5215 0.8155 0.7020 MGDN 0.7013 0.4222 0.8355 23.7661 0.6773 0.6331
PADCDTNP 0.7613 0.5916 0.8417 26.9731 0.8394 0.8075 PADCDTNP 0.7367 0.5268 0.8224 23.1335 0.7104 0.7274
DB-MFIF 0.7479 0.5027 0.8402 26.5203 0.8398 0.7775 DB-MFIF 0.6958 0.4290 0.8207 23.3339 0.6620 0.6642
@@
SAMF 0.7511 0.5585 0.8401 26.8899 0.8256 0.7951 SAMF 0.6292 0.3009 0.7984 23.2530 0.5586 0.6693
TC-MoA 0.7401 0.5274 0.8415 26.1720 0.8188 0.7597 TC-MoA 0.6034 0.2826 0.8216 23.4429 0.5307 0.6475
Mask-DiFuser 0.6004 0.3917 0.8182 25.0254 0.6765 0.5991 Mask-DiFuser 0.5585 0.3556 0.7889 22.6972 0.5367 0.5486
DMANet 0.7606 0.5807 0.8396 26.9807 0.8439 0.8055 DMANet 0.7384 0.5247 0.8119 23.6877 0.7073 0.7279
GIFNet 0.5194 0.3268 0.7854 25.9368 0.5412 0.6042 GIFNet 0.4509 0.2524 0.7653 22.5889 0.3850 0.5433
@@
MGDN 0.7230 0.5160 0.8389 27.1386 0.7527 0.7570 MGDN 0.8042 0.5050 0.9552 33.3076 0.8902 0.7849
PADCDTNP 0.7311 0.6193 0.8316 27.0261 0.7526 0.8231 PADCDTNP 0.8203 0.6502 0.9517 32.1256 0.9213 0.8318
DB-MFIF 0.7192 0.5278 0.8317 27.6825 0.7509 0.7813 DB-MFIF 0.7934 0.5353 0.9481 33.4508 0.8940 0.7990
@@
SAMF 0.7263 0.6248 0.8304 27.6122 0.7571 0.8203 SAMF 0.7093 0.3293 0.9330 30.8761 0.7215 0.7707
TC-MoA 0.6557 0.5170 0.8388 27.6032 0.7098 0.7660 TC-MoA 0.7915 0.5109 0.9501 32.8831 0.8906 0.7971
Mask-DiFuser 0.5573 0.4306 0.8031 26.3362 0.6230 0.6032 Mask-DiFuser 0.6561 0.4534 0.8446 31.5767 0.7927 0.6520
DMANet 0.7315 0.6248 0.8302 27.6611 0.7567 0.8237 DMANet 0.8205 0.6484 0.9507 34.0174 0.9194 0.8362
GIFNet 0.4388 0.3536 0.7523 26.1136 0.4642 0.5500 GIFNet 0.4834 0.3572 0.8471 30.5346 0.6467 0.6041

### v3

@@ Abstract
Multi-focus image fusion (MFIF) is a crucial technique in image processing, with a key challenge being the generation of decision maps with precise boundaries. However, traditional methods based on heuristic rules and deep learning methods with black-box networks are difficult to generate high-quality decision maps. To overcome this challenge, we introduce neurodynamics-driven coupled neural P (CNP) systems, which are biological neural computation models inspired by spiking mechanisms, to enhance the accuracy of decision maps. Specifically, we first conduct an in-depth analysis of the model’s neurodynamics to identify the constraints between the network parameters and the input signals. This solid analysis avoids abnormal continuous firing of neurons and ensures the model accurately distinguishes between focused and unfocused regions, generating high-quality decision maps for MFIF. Based on this analysis, we propose a Neurodynamics-Driven CNP Fusion model (ND-CNPFuse) tailored for the challenging MFIF task. Unlike current ideas of decision map generation, ND-CNPFuse distinguishes between focused and unfocused regions by mapping the source image into interpretable spike matrices. By comparing the number of spikes, an accurate decision map can be generated directly without any post-processing. Extensive experimental results show that ND-CNPFuse achieves new state-of-the-art performance on four classical MFIF datasets, including Lytro, MFFW, MFI-WHU, and Real-MFF. The code is available at https://github.com/MorvanLi/ND-CNPFuse.
@@ 1 Introduction
Multi-focus image fusion (MFIF) [45] is an important task in computer vision that aims to generate an all-in-focus image by fusing multiple images of the same scene captured at varying focal depths. The fused image improves both perceptual quality and information richness, benefiting downstream vision applications such as action detection [32] and object detection [17]. By integrating the focused regions from source images with different focus settings, MFIF effectively alleviates the blurring artifacts introduced by the limited depth of field inherent in imaging systems.
@@
Figure 1: Workflow comparison of the existing decision method and the proposed ND-CNPFuse. ND-CNPFuse encodes source images into spike matrices and generates the decision map by comparing spike counts. Focused regions produce more spikes than unfocused ones, consistent with human visual perception [12].
@@
Before the advent of deep learning (DL), MFIF methods were performed in the transform [19] and spatial domains [46]. These methods often rely heavily on manually designed fusion rules and strong prior assumptions, which may not always be applicable or sufficiently accurate.
@@
In recent years, DL-based approaches have become the dominant paradigm in MFIF, achieving remarkable progress [22]. Some studies have applied DL techniques to end-to-end image fusion frameworks [41, 37, 30, 14, 7]. Different end-to-end methods may use various network structures, but they generally consist of an encoder network for extracting image features, a fusion network for merging features, and a decoder network for reconstructing the fused features. Another line of studies draws inspiration from spatial domain methods, suggesting the use of DL techniques to generate decision maps [23, 33, 28, 48, 10]. This process involves training a network to map the source image to a focus map, ultimately producing decision maps. Compared to end-to-end methods, decision-based methods directly retain information from the source image at the pixel level, avoiding the spatial consistency issues of end-to-end methods. Therefore, more research has been done on DL-based decision map methods. However, DL-based decision map methods still have evident drawbacks. The most crucial one is that the internal working mechanism is difficult to explain, hindering a clear understanding of the mapping between source images and the decision map [34]. This will result in nuisance false edges and burrs in the generated decision map, as can be intuitively observed in Fig. 1(a).
@@
Coupled neural P (CNP) systems [27] are a class of neural-like computing models inspired by synchronous pulse bursts in a mammal’s visual cortex. A CNP system generates synchronous spikes through the mutual coupling of neurons, where synchronous pulses of neurons are treated as 0 or 1. After multiple iterations, the system produces a sequence of binary images that record the firing states of each neuron at every time step. By comparing the spike counts of neurons, the degree of focus corresponding to each pixel can be effectively estimated, enabling the differentiation between focused and defocused regions. This is particularly well-suited for decision map generation in the MFIF task, which serves as the core motivation for introducing CNP systems into MFIF. However, when CNP systems are directly applied to MFIF, neurons may suffer abnormal continuous firing. In this case, the spike counts cannot accurately reflect the focus differences of the source images, which affects the quality of the decision map. This issue stems from an inappropriate constraint between network parameters and input signals.
@@
To address this problem, we first conduct a detailed study on the neurodynamics of CNP neurons. This analysis reveals the underlying constraints between network parameters and input signals, thereby preventing abnormal continuous firing in neurons. A solid understanding of the neural firing mechanism is helpful for setting appropriate parameters and provides a theoretical basis for effectively applying CNP systems in MFIF. Building on the neurodynamics-driven CNP system, we then propose a new fusion method called ND-CNPFuse. As illustrated in Fig. 1(b), ND-CNPFuse provides an interpretable and accurate decision map generation workflow, in contrast to existing DL-based MFIF methods that rely on black-box networks (Fig. 1(a)).
@@ 2.1 DL-based MFIF methods
End-to-end methods. These methods directly generate fused images by leveraging the fitting ability of deep neural networks. Representative approaches include convolutional neural networks [47, 36, 42, 4], generative adversarial networks [40, 8], Transformer architectures [24, 49], and diffusion models [2, 31, 39]. However, maintaining spatial consistency with source images remains challenging.
Decision-map methods. These methods formulate MFIF as a pixel-level classification task by predicting decision maps for focused and defocused regions. Liu et al. [18] first applied this technique to MFIF, but the generated decision maps remain limited in precision. Subsequent studies improved performance using ensemble learning [1], attention mechanisms [35], regression pair learning [16], multi-scale architectures [21], edge preservation [34], and defocus blur modeling [29]. Nevertheless, unclear network mechanisms often introduce pseudo-edges and burr artifacts, and generating high-quality decision maps remains to be explored.
@@ 3.1 Mathematical model of CNP systems
$T_{ij}(t)=\left\{\begin{array}[]{ll}T_{ij}(t-1)-\tau+\lambda P_{ij}(t-1),&\text{if}~\sigma_{ij}~\text{fires}\\ T_{ij}(t-1),&\text{otherwise}\end{array}\right$. (5)
where subscripts $\left(i,j\right)$ denote the coordinates of the neuron, $\left(k,l\right)$ denote the coordinates of its neighboring neurons, and $t$ represents the current iteration count. $u,v,\tau$ indicate the spikes consumed by three units. $I_{ij}$ is an external input. $W_{kl}$ characterizes synaptic weight. $P_{kl}$ symbolizes spikes received from the surrounding neurons $\sigma_{kl}$, where neighborhood radius is $r$ ($\sigma_{kl}\in\delta_{r}$). $\lambda$ is a threshold weight.
@@ 3.2 Why neurodynamics analysis is necessary?
In certain cases, CNP neurons may enter a state of continuous firing, generating spikes at each time step. As illustrated in Fig. 2(c) and Fig. 2(d). This uncontrolled spiking behavior can lead to model failure, as the output no longer reflects the features of the input signal. For the MFIF task, such outputs do not effectively distinguish between focused and defocused regions, thereby affecting the generation of accurate decision maps. We next analyze this invalid state and identify parameter configurations that prevent it.
Analysis of continuous firing condition. We analyze the continuous firing condition of CNP neurons to derive constraints between parameters and the input signal. For continuously firing CNP neurons, the state update is given by:
$U(t)$ $=U(t-1)-u+I+\sum\limits_{\sigma_{kl}\in\delta_{r}}W_{kl}P_{kl}(t-1),$ (6)
$V(t)$ $=V(t-1)-v+\sum\limits_{\sigma_{kl}\in\delta_{r}}W_{kl}P_{kl}(t-1),$ (7)
$T(t)$ $=T(t-1)-\tau+\lambda P(t-1).$ (8)
Following [27], we approximate the spike consumption of the three CNP units as linear decay to facilitate analysis.
@@
$U(t)$ $=I\frac{1-\alpha^{t}}{1-\alpha}+\sum_{n=0}^{t-1}K(n)\alpha^{t-n-1}.$ (10)
@@
Let us clarify the initial conditions: $U(0)$ and $K(0)$ are both zero. Define $K(n)=\sum_{\sigma_{kl}\in\delta_{r}}W_{kl}P_{kl}(n)$. Next, we describe $U(1),U(2),U(3),\ldots,U(t)$ step by step:
@@
$U(t)$ $=I\frac{1-\alpha^{t}}{1-\alpha}+\sum_{n=0}^{t-1}K(n)\alpha^{t-n-1}.$
@@
[…]
@@
$\begin{split}&\biggl(I\frac{1-\alpha^{t}}{1-\alpha}+\sum_{n=0}^{t-1}K(n)\alpha^{t-n-1}\biggl)\biggl(1+\sum_{n=0}^{t-1}K(n)\beta^{t-n-1}\biggl)\\ &\quad\quad\quad\quad\quad\quad\quad\quad>\lambda\frac{1-\gamma^{t-1}}{1-\gamma}.\end{split}$ (17)
Since the central CNP neuron and its surrounding neurons all continue to fire, that is, $P_{kl}(t-1)\equiv 1$, and we obtain $K(n)=\mathrm{sum}(W)$, where $\mathrm{sum}(W)$ denotes the sum of its elements. Therefore, Eq. (17) can be expressed as:
@@
When $t$ is sufficiently large, the system approaches stability. We can obtain Eq. (19):
@@
According to Theorem 4, the external input to a CNP neuron should remain below its internal state to prevent uncontrolled continuous firing, as formulated below
$I\leq\frac{\lambda(1-\alpha)(1-\beta)}{(1-\gamma)(1-\beta+\operatorname{sum}(W))}-\operatorname{sum}(W).$ (21)
Corollary 1 reveals the necessary constraints between parameters and input signals for the proper functioning of CNP neurons. Detailed configurations of CNP neuron parameters such as $u$, $v$, and $\tau$ derived from Corollary 1 are provided in Appendix A.2. Noting that the internal parameters of neurons are automatically configured based on input MFIF images, eliminating the manual tuning, and demonstrating transferability and generalization across different datasets. This helps to improve the ability of CNP systems to distinguish between focused and defocused regions in MFIF tasks, thereby improving the quality of the generated decision map.
@@ 4.2 Overview of the architecture
CCF [NeurIPS2024] 0.4907 0.3426 0.8525 25.1774 0.5868 0.6121 0.4496 0.2866 0.8384 22.6547 0.4320 0.5498
PADCDTNP [NN2024] 0.7613 0.5916 0.8417 26.9731 0.8394 0.8075 0.7367 0.5268 0.8224 23.1335 0.7104 0.7274
DB-MFIF [TMM2024] 0.7479 0.5027 0.8402 26.5203 0.8398 0.7775 0.6958 0.4290 0.8207 23.3339 0.6620 0.6642
@@
SAMF [ICASSP2024] 0.7511 0.5585 0.8401 26.8899 0.8256 0.7951 0.6292 0.3009 0.7984 23.2530 0.5586 0.6693
TC-MoA [CVPR2024] 0.7401 0.5274 0.8415 26.1720 0.8188 0.7597 0.6034 0.2826 0.8216 23.4429 0.5307 0.6475
Mask-DiFuser [TPAMI2025] 0.6004 0.3917 0.8182 25.0254 0.6765 0.5991 0.5585 0.3556 0.7889 22.6972 0.5367 0.5486
DMANet [AAAI2025] 0.7606 0.5807 0.8396 26.9807 0.8439 0.8055 0.7384 0.5247 0.8119 23.6877 0.7073 0.7279
GIFNet [CVPR2025] 0.5194 0.3268 0.7854 25.9368 0.5412 0.6042 0.4509 0.2524 0.7653 22.5889 0.3850 0.5433
@@
CCF [NeurIPS2024] 0.4972 0.3793 0.8389 25.9154 0.5614 0.6691 0.7179 0.4120 0.9458 30.0386 0.7743 0.7118
PADCDTNP [NN2024] 0.7311 0.6193 0.8316 27.0261 0.7526 0.8231 0.8203 0.6502 0.9517 32.1256 0.9213 0.8318
DB-MFIF [TMM2024] 0.7192 0.5278 0.8317 27.6825 0.7509 0.7813 0.7934 0.5353 0.9481 33.4508 0.8940 0.7990
@@
SAMF [ICASSP2024] 0.7263 0.6248 0.8304 27.6122 0.7571 0.8203 0.7093 0.3293 0.9330 30.8761 0.7215 0.7707
TC-MoA [CVPR2024] 0.6557 0.5170 0.8388 27.6032 0.7098 0.7660 0.7915 0.5109 0.9501 32.8831 0.8906 0.7971
Mask-DiFuser [TPAMI2025] 0.5573 0.4306 0.8031 26.3362 0.6230 0.6032 0.6561 0.4534 0.8446 31.5767 0.7927 0.6520
DMANet [AAAI2025] 0.7315 0.6248 0.8302 27.6611 0.7567 0.8237 0.8205 0.6484 0.9507 34.0174 0.9194 0.8362
GIFNet [CVPR2025] 0.4388 0.3536 0.7523 26.1136 0.4642 0.5500 0.4834 0.3572 0.8471 30.5346 0.6467 0.6041